# m1: Unleash the Potential of Test-Time Scaling for Medical Reasoning with Large Language Models

### v1

@@ 5 Conclusions
We introduced m1, demonstrating that test-time scaling significantly improves medical reasoning in large language models without requiring extensive fine-tuning. Performance across diverse medical QA benchmarks consistently improved with increased inference-time reasoning budgets. Crucially, we found that test-time scaling alone cannot remedy fundamental deficiencies in medical knowledge, emphasizing the necessity for high-quality medical data and model scale expansion. m1 achieves strong performance, outperforming more expensive approaches such as HuatuoGPT-o1 and UltraMedical on various benchmarks. Our 7B model trained on 23K data establishes a new state-of-the-art in the $\leq 10B$ parameter category, and our 32B model rivals models 2×2$\times$2 × in size. We release a full-stack open-source package including the curated dataset (m1K), fine-tuned model weights, and inference code with budget control, to encourage future exploration in optimizing inference strategies in clinical AI applications.
@@ Appendix A Data Statistics
As illustrated in Figure 6, the distributions of training token lengths between the m1K/m23K and s1K (Muennighoff et al., 2025) datasets exhibit clear differences. The m1K/m23K dataset shows a strongly right-skewed distribution, with most samples having token lengths clustered around 1,000 tokens, quickly diminishing toward lengths beyond 3,000 tokens. In contrast, the s1K dataset displays a more uniform and broader distribution, spanning widely from about 2,500 to over 15,000 tokens, with peaks around 5,000 to 10,000 tokens. These contrasting distributions reflect differing data preparation strategies: m1K/m23K focuses on concise medical knowledge without longer thinking steps, whereas s1K includes longer, more detailed reasoning traces suitable for complex multi-step inference tasks.
(a) m1K
(b) m23K
(c) s1K
Figure 6: The token length distributions of m1K and s1K (Muennighoff et al., 2025). The 25%/50%/75% quantile is marked in transparent vertical dotted lines.
@@ B.2 SFT Details
All fine-tuning runs use standard language modeling training with ‘trl‘ library: we optimize the model to minimize the output cross-entropy on the reasoning $\to$ answer sequences (teacher-forcing the entire sequence). We use the same training hyperparameters as the s1 paper for consistency (Muennighoff et al., 2025): 5 epochs of training, a batch size equals 16, a low learning rate of 1e-4 with warmup and cosine decay, a modest weight decay of 1e-4, Adam betas of 0.9 and 0,95. The thinking part is enclosed with <|im_start|>think and <|im_start|>answer. SFT is performed with trl and transformers libraries.
@@ B.3 Evaluation Details
    MedMCQA (Pal et al., 2022) – a collection of  3.5K multiple-choice questions from Indian medical entrance exams, testing general medical knowledge.
@@
    MedQA-USMLE (Jin et al., 2021) – the USMLE question dataset (NYU MedQA) containing US medical licensing exam MCQs; we use the standard test split.
@@
    PubMedQA (Jin et al., 2019) – a dataset of biomedical research questions (factoid Q paired with abstracts) where the task is to answer yes/no/maybe or short answer. These three were part of our training data pool (though we filtered and sampled from them), so they represent in-domain evaluations. We report accuracy (for MCQ, percentage of correct choices; for PubMedQA, percentage of correct yes/no/maybe).
@@
    MMLU-Pro (Wang et al., 2024) (Medical) – the medical category subset of the Massive Multitask Language Understanding benchmark, which includes professional medicine questions and related subjects. We specifically evaluate on the Professional Medicine section (and report accuracy). We follow the split from (Chen et al., 2024b).
@@
    GPQA (Medical) (Rein et al., 2024) – the biomedical portion of the Graduate-Level Physics/Chemistry/Biology QA dataset GPQA. This dataset contains extremely challenging, Google-proof multiple-choice questions created by experts, requiring high reasoning (we use the biology/medical questions,  150 in total). We follow the split from (Chen et al., 2024b).
@@
    MedBullets (Chen et al., 2024a) – a collection of practice questions from the MedBullets medical education platform. We specifically take subsets of difficulty level 4 and 5 (on a 1–5 scale, 5 being hardest), denoted MedBullets Op4 and MedBullets Op5, about 100 questions each, to serve as challenging test sets.
@@
    MedXpertQA (Zuo et al., 2025) – a custom set of 50 expert-written multi-step medical reasoning questions we created for qualitative evaluation (free-form answers). For MedXpertQA we report the percentage of questions answered correctly.
@@ Methods.
    OpenBioLLM (8B) (Pal & Sankarasubbu, 2024) – The 8B model from Saama AI, fine-tuned with expert-curated medical data.
@@
    MMed-Llama (8B) (Qiu et al., 2024) – A multilingual medical model from MedS3 work, which underwent additional pre-training (denoted MMedS or MMed in results).
@@
    Med42 (8B) (Christophe et al., 2024) – The 8B model from the Med42-v2 suite, instruction and preference-tuned on clinical data.
@@
    UltraMedical (8B) (Zhang et al., 2024) – The 8B model from Tsinghua’s UltraMedical project (we test both the v3.0 and v3.1 versions if available).
@@
    HuatuoGPT-o1 (7B & 8B) (Chen et al., 2024b) – The smaller versions of HuatuoGPT-o1 (the 70B model’s distilled or intermediate checkpoints) as reported in their paper.
@@
    Larger models ($>$10B): We also compare to state-of-the-art open models in the larger size class: Med42-70B (Christophe et al., 2024), OpenBioLLM-70B (Pal & Sankarasubbu, 2024), UltraMedical-70B (Zhang et al., 2024), and HuatuoGPT-o1-70B/72B (Chen et al., 2024b) (if available). These represent the current best open medical LLMs (some claim parity with GPT-4).
It is worth noting that some of these baselines (e.g. HuatuoGPT-o1 (Chen et al., 2024b), UltraMedical  (Zhang et al., 2024)) involve complex training regimes (RL or extensive preference tuning), and in cases like Med42 and OpenBioLLM, they incorporate expert feedback. Our approach does not, so beating or matching them would be a strong indication of the power of test-time scaling.

### v2

@@ 5 Conclusions
We introduced m1, demonstrating that test-time scaling significantly improves medical reasoning in large language models without requiring extensive fine-tuning. Performance across diverse medical QA benchmarks consistently improved with increased inference-time reasoning budgets. Crucially, we found that test-time scaling alone cannot remedy fundamental deficiencies in medical knowledge, emphasizing the necessity for high-quality medical data and model scale expansion. m1 achieves strong performance, outperforming more expensive approaches such as HuatuoGPT-o1 and UltraMedical on various benchmarks. Our 7B model trained on 23K data establishes a new state-of-the-art in the $\leq 10B$ parameter category, and our 32B model rivals models $2\times$ in size. We release a full-stack open-source package including the curated dataset (m1K), fine-tuned model weights, and inference code with budget control, to encourage future exploration in optimizing inference strategies in clinical AI applications.
@@ Appendix A Data Statistics
As illustrated in Figure 8(a), the distributions of training token lengths between the m1K/m23K and s1K muennighoff2025s1 datasets exhibit clear differences. The m1K/m23K dataset shows a strongly right-skewed distribution, with most samples having token lengths clustered around 1,000 tokens, quickly diminishing toward lengths beyond 3,000 tokens. In contrast, the s1K dataset displays a more uniform and broader distribution, spanning widely from about 2,500 to over 15,000 tokens, with peaks around 5,000 to 10,000 tokens. These contrasting distributions reflect differing data preparation strategies: m1K/m23K focuses on concise medical knowledge without longer thinking steps, whereas s1K includes longer, more detailed reasoning traces suitable for complex multi-step inference tasks.
(a) The token length distributions of m1K and s1K muennighoff2025s1. The 25%/50%/75% quantile is marked in transparent vertical dotted lines.
@@ B.2 SFT Details
All fine-tuning runs use standard language modeling training with ‘trl‘ library: we optimize the model to minimize the output cross-entropy on the reasoning $\to$ answer sequences (teacher-forcing the entire sequence). We use the same training hyperparameters as the s1 paper for consistency muennighoff2025s1: 5 epochs of training, a batch size equals 16, a low learning rate of 1e-4 with warmup and cosine decay, a modest weight decay of 1e-4, Adam betas of 0.9 and 0.95. The thinking part is enclosed with <|im_start|>think and <|im_start|>answer. SFT is performed with trl and transformers libraries.
@@ B.3 Evaluation Details
    MedMCQA pal2022medmcqa – a collection of  3.5K multiple-choice questions from Indian medical entrance exams, testing general medical knowledge.
@@
    MedQA-USMLE jin2021disease – the USMLE question dataset (NYU MedQA) containing US medical licensing exam MCQs; we use the standard test split.
@@
    PubMedQA jin2019pubmedqa – a dataset of biomedical research questions (factoid Q paired with abstracts) where the task is to answer yes/no/maybe or short answer. These three were part of our training data pool (though we filtered and sampled from them), so they represent in-domain evaluations. We report accuracy (for MCQ, percentage of correct choices; for PubMedQA, percentage of correct yes/no/maybe).
@@
    MMLU-Pro wang2024mmlu (Medical) – the medical category subset of the Massive Multitask Language Understanding benchmark, which includes professional medicine questions and related subjects. We specifically evaluate on the Professional Medicine section (and report accuracy). We follow the split from chen2024huatuogpt.
@@
    GPQA (Medical) rein2024gpqa – the biomedical portion of the Graduate-Level Physics/Chemistry/Biology QA dataset GPQA. This dataset contains extremely challenging, Google-proof multiple-choice questions created by experts, requiring high reasoning (we use the biology/medical questions,  150 in total). We follow the split from chen2024huatuogpt.
@@
    MedBullets chen2024benchmarking – a collection of practice questions from the MedBullets medical education platform. We specifically take subsets of difficulty level 4 and 5 (on a 1–5 scale, 5 being hardest), denoted MedBullets Op4 and MedBullets Op5, about 100 questions each, to serve as challenging test sets.
@@
    MedXpertQA zuo2025medxpertqa – a custom set of 50 expert-written multi-step medical reasoning questions we created for qualitative evaluation (free-form answers). For MedXpertQA we report the percentage of questions answered correctly.
@@ Methods.
    OpenBioLLM (8B) pal2024openbiollms – The 8B model from Saama AI, fine-tuned with expert-curated medical data.
@@
    MMed-Llama (8B) qiu2024towards – A multilingual medical model from MedS3 work, which underwent additional pre-training (denoted MMedS or MMed in results).
@@
    Med42 (8B) christophe2024med42 – The 8B model from the Med42-v2 suite, instruction and preference-tuned on clinical data.
@@
    UltraMedical (8B) zhang2024ultramedical – The 8B model from Tsinghua’s UltraMedical project (we test both the v3.0 and v3.1 versions if available).
@@
    HuatuoGPT-o1 (7B & 8B) chen2024huatuogpt – The smaller versions of HuatuoGPT-o1 (the 70B model’s distilled or intermediate checkpoints) as reported in their paper.
@@
    Larger models (>>10B): We also compare to state-of-the-art open models in the larger size class: Med42-70B christophe2024med42, OpenBioLLM-70B pal2024openbiollms, UltraMedical-70B zhang2024ultramedical, and HuatuoGPT-o1-70B/72B chen2024huatuogpt (if available). These represent the current best open medical LLMs (some claim parity with GPT-4).
It is worth noting that some of these baselines (e.g. HuatuoGPT-o1 chen2024huatuogpt, UltraMedical  zhang2024ultramedical) involve complex training regimes (RL or extensive preference tuning), and in cases like Med42 and OpenBioLLM, they incorporate expert feedback. Our approach does not, so beating or matching them would be a strong indication of the power of test-time scaling.